OpenStudy (anonymous):

Find the probability. A die is rolled 20 times and the number of twos that come up is tallied. Find the probability of getting exactly five twos. A. 0.129 B. 0.921 C. 0.003 D. 0.083

OpenStudy (anonymous):

exactly 5 twos for this binomial probability probability you get a 2 on any one toss is \(\frac{1}{6}\) probability you do not is \(\frac{5}{6}\) use \[\dbinom{20}{5}\times (\frac{1}{6})^5\times (\frac{5}{6})^{15}\] and a calculator

OpenStudy (anonymous):

can you just tell me the answer ?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

Math Is so hard for me .